OpenStudy (anonymous):

Maria received a gift of cash from her grandmother. She put two-thirds of the money in her savings account, spent twenty-five dollars on a pair of jeans, and spent half the remaining money on a DVD. If Maria has eighteen dollars left over, how much money did she receive from her grandmother?

OpenStudy (anonymous):

She got a lot of money... I'm kidding here's how you find it Work backwards She spent half of the remaining money on a DVD If she has 18 dollars left...what is 18*2?

OpenStudy (anonymous):

I was going to give a full walkthrough but since you're not replying, ill let the other guy take over

OpenStudy (whpalmer4):

If for some reason you don't feel comfortable working backwards, and want to do it by solving an equation: Let \(x\) be the amount she is given. \[x-\frac{2}{3}x= \text{amount she has left after depositing 2/3 in savings account}\]\[x-\frac{2}{3}x - 25=\text{amount she has left after buying a pair of jeans for \$25}\]\[(\frac{1}{2})(x-\frac{2}{3}x-25)=\text{amount she has left after buying DVD}\] and you know that final amount is \($18\) so your final equation before simplification is\[18 = (\frac{1}{2})(x-\frac{2}{3}x-25)\] Pretty easy to solve that for \(x\). Be sure to check your answer, whichever approach you take to finding it!